# AI Assistants to Enhance and Exploit

### v1

@@ Abstract
I Introduction
High-performance scientific computing—which underpins discovery and innovation in all areas of science, engineering, technology, and society—relies on complex, evolving software libraries and tools that demand deep domain knowledge to develop, use, and maintain across continual advances in science drivers and computing architectures. Generative AI provides unprecedented opportunities to transform strategies for research and development (R&D) of scientific software, with the ultimate goal of accelerating scientific insight [1, 2, 3]. PETSc—the Portable, Extensible Toolkit for Scientific Computation—has served the scientific computing community for over three decades as a foundational software infrastructure for scalable numerical solvers [4, 5, 6]. During this time, the project has cultivated a sophisticated and robust code base and a deep, distributed body of developer and user knowledge. The community knowledge base includes source code and documentation, email exchanges in public mailing lists, GitLab issues, git merge request reviews, Discord chats, manual pages and examples from the PETSc website, written materials such as papers and presentations, and importantly, PETSc developers and users’ understanding of PETSc. Among the written material, the structured and curated subset constitutes the official knowledge base: documentation, code, frequently asked questions (FAQs), and other resources that have undergone refinement and review, essentially everything that has passed a git merge request, while the remaining is the unofficial knowledge base. We call the subset of community knowledge that is available only in a developer’s brain (i.e., not written in a machine-accessible format) the wet knowledge base [^1] https://en.wikipedia.org/wiki/Wetware_(brain), and the opposite the dry knowledge base, see Fig. 1.
Figure 1: PETSc community knowledge base. AI assistants could help the flow of knowledge in all directions.
@@ I Introduction
These tools do not replace human contributors but act as scalable collaborators—suggesting code, summarizing discussions, generating documentation updates, and responding to user questions while adhering to PETSc’s collaborative development process. These same tools and techniques can and will be used to transfer material from the unofficial knowledge base to the official knowledge base, for example, dramatically improving manual pages with little developer effort. Developer-user (as well as developer-developer) communication is a common way that information is transferred from the wet knowledge base to machine-accessible format. Questions provoke developers to write down (often informally) knowledge and conceptualizations that they have never written down before. By having the LLM process these developer-user interactions, we can effectively capture wet knowledge as the developers perform their daily tasks without requiring them to try to core-dump their wet knowledge actively.
@@
Figure 2: Our goal is to create PETSc AI assistants capable of understanding and executing scientific computing tasks at the heart of numerical library research, providing the AI equivalent of human PETSc developers using LLM technology, while also continually advancing the PETSc knowledge base.
@@
Assistants embedded in tools such as Discord can answer user queries, recommend solutions, aid onboarding, and even analyze recurring feedback to guide improvements. For PETSc, these capabilities help address questions on the development of applications that build on PETSc scalable linear, nonlinear, and timestepping solvers as well as lower-level vector and matrix data structures for heterogeneous computing architectures [7] and outer-loop numerical optimization and analytics.
@@
AI assistants can help in code review, suggest code improvements, aid in API design reviews, offer performance tuning strategies, assist with refactoring complex code, and develop completely new functionality. Such assistants could help with continual advances needed in the PETSc code base to achieve efficient performance and address next-generation science challenges on ever-changing computer architectures.
@@
LLMs can help generate automated tests, detect bugs, and assist in debugging or improving continuous integration practices. Such capabilities would help open-source libraries such as PETSc achieve comprehensive, multilevel testing to ensure a robust and reliable code base as part of the broader scientific software ecosystem.
@@
Beyond software development and maintenance, LLMs can support scientific discovery itself—by aiding in literature reviews [8], surfacing unexplored research directions [9], and suggesting algorithms for novel problems. Research in numerical libraries incorporates advances in algorithms and software, as driven by the needs of next-generation scientific applications and the realities of continually changing computer architectures.
@@
Through meeting summarization, task prioritization, and resource tracking, AI tools can reduce cognitive load and coordination overhead among contributors. For PETSc developers, such capabilities help with communication and synchronization across a multi-institutional distributed team.
@@
II PETSc User Support
User support, which we have consistently provided for over thirty years, is a fundamental responsibility of the PETSc team—as it is for any open-source scientific software project—to disseminate knowledge and realize the full impact of the software. Consequently, user support is our first focus when exploiting the PETSc knowledge base with AI assistants. Like many scientific computing libraries, PETSc has been funded primarily by the U.S. government. However, not only has this funding been limited, but it has also largely focused on supporting research in numerical computing, leaving few resources available for software maintenance and user support. Over the years, PETSc has developed a large and diverse user base, with tens of thousands of users across dozens of scientific domains. As a result, user support has become a significant undertaking. For instance, in May 2025 alone, there were over 300 messages across PETSc mailing lists, Discord channels, and GitLab issues. While PETSc developers have remained deeply committed to supporting users, the volume and complexity of user inquiries can often overwhelm a small team, particularly when resolving issues that involve extended, multiday, or multiweek discussions. Thus, to ensure that this process remains sustainable in the long term, we are particularly interested in leveraging recent advances in AI to assist with user support.
Although PETSc users can employ general-purpose AI tools such as ChatGPT, these tools can produce incorrect or hallucinated responses when answering technical questions about PETSc. The reason is primarily that mainstream LLMs do not have access to sufficient PETSc information. For example, in January 2025, we posed a PETSc-specific question to a paid version of ChatGPT and observed such limitations firsthand.
@@ II PETSc User Support
We scored this response as 0: it was a hallucination to a question about “KSPBurb” – a fictitious term that follows the naming convention of PETSc Krylov solvers (see Table I for the rubric used in the assessment). Although later versions of ChatGPT gave better results, this example shows the risk of relying solely on mainstream LLMs. The example also reminds us of the dangers of incorporating AI into user support. As noted in [10], a prominent platform experienced backlash when its AI agents, masquerading as human support staff, fabricated a nonexistent login policy in response to user inquiries that led to widespread account cancellations. To safeguard the integrity and reputation of PETSc, we initially have adopted a cautious approach, and to date we have avoided full automation in user support interactions.
@@
To address the above challenges, we propose in Section III an augmented PETSc LLM workflow that can effectively process and deliver PETSc-centric information, and in Section IV we present an integrated LLMs and PETSc Discord bots system that can reduce the risk of hallucinated responses reaching users.
III The PETSc Augmented LLM Workflow
We have developed a preliminary PETSc LLM workflow, illustrated in Fig. 3, which can be used in multiple settings. For example, we could use the LLM to answer questions from the PETSc Discord server, from mailing lists, or from a chat box on the PETSc homepage. For developers, we could even provide command line tools and integrated development environment (IDE) extensions to facilitate various use cases. Our approach builds on standard LLM tools, such as LangChain [11], while incorporating practical adaptations for PETSc. Specifically, we do the following:
@@ III The PETSc Augmented LLM Workflow
Figure 3: Our approach for creating PETSc AI assistants builds on standard LLM ecosystem tools. The boxes 1, 2, 3, and 4 in the top bar indicate the sequence of processing for PETSc-specific LLM interactions.
III-A Generating the RAG databases
Retrieval-augmented generation algorithms ensure that users receive the most contextually appropriate and accurate information. RAG consists of two phases: the generation phase and the retrieval phase. Here, we discuss the generation phase. PETSc documentation is maintained in Markdown format and processed via Sphinx [12] into the website format. To construct a practical RAG database using this documentation, we use the following steps (see Fig. 3, beginning with “PETSc knowledge base”, following the blue arrows to “Embedding generator” and then the green arrow to “RAG databases”):
@@ III-A Generating the RAG databases
III-B Utilizing the documentation RAG database
@@
III-C Augmenting RAG searches with PETSc-specific keyword searches
@@ III-D Reranking enhanced RAG
Figure 4: Reranking-enhanced RAG workflow.
III-E Postprocessing LLM output for users
@@
III-F Storing and tracking interaction history
In addition to visually presenting the LLM output, we keep a detailed, manipulatable, searchable database of all interactions with all the LLMs by the PETSc developers (and eventually also users). In Fig. 3 this is depicted with the dotted orange arrow from box 4 to “Shared histories”; the blue dotted line to box 1 indicates that material from the shared history will also eventually be included in the RAG and reranking processing and passed to the LLM. Questions, responses, dates when done, and the continuation and embedding model used for each interaction will be included, as well as the generated prompts in the interaction. We provide an easy way for “scorers” to “blind-score” the responses to questions and indicate correct and incorrect portions of the responses. We can also score answers from PETSc developers stored in the same database, establishing high quality and trust. Currently, we use a bespoke Python dictionary to manage the database, but we are pursuing using emerging agentic memory systems [13].
IV PETSc AI Bots in Discord
Discord [14] is an instant messaging platform where communication happens in virtual communities called “servers.” We have set up a PETSc Discord server, and hundreds of users have joined since its debut. We integrated augmented PETSc LLMs with the Discord server and set up a chatbot that has two usage modes: (1) one can directly chat with it, or (2) the chatbot can answer questions from emails, while PETSc developers can monitor and vet answers on back channels. In this section we first briefly introduce Discord’s integration mechanisms, then describe the workflow that connects a PETSc public mailing list and LLMs; see Fig. 5.
Figure 5: Bots in Discord that connect a PETSc mailing list and an LLM. Numbered arcs show a typical sequence of events in the workflow.
@@ IV PETSc AI Bots in Discord
PETSc has three public mailing lists: {petsc-users, petsc-maint, petsc-dev}. petsc-dev is mainly for PETSc developers and others interested in PETSc development. The other two are for PETSc users with one distinction: petsc-maint is a private maintenance email with controlled recipients and no public archives, while petsc-users is fully public and has archives from the past twenty years. In this study we targeted petsc-users but didn’t touch its archives for RAG. We created a Google Gmail account petscbot@gmail.com and subscribed it to the petsc-users mailing list. Then, with Google Apps Script services [15], we use JavaScript to periodically check whether there are new (unread) emails from petsc-users in the Gmail account. If there are, the script sends a message to a webhook associated with a private channel named petsc-users-notification in PETSc Discord. An email bot (application) was added to the PETSc Discord server, which watches for new messages in petsc-users-notification. Whenever there is a new message, using Gmail’s Python API, the email bot fetches unread emails from petsc-users in the Gmail account and marks them read. The fetched emails are posted to a private channel, visible only to PETSc developers, named petsc-users-emails. This channel is a Discord Forum channel made of posts. We can make each email thread a post, with its title being the email thread’s subject and emails within the thread follow up messages in the post. We also fetch and attach email attachments to their corresponding messages in Discord. We lightly parse email bodies to remove quotes commonly seen in email replies and revert the url-defense protected URLs so that messages are presented concisely.
@@
From a post in petsc-users-emails, any PETSc developer can invoke a command /reply on the chatbot, which will then build a conversation context with the title, messages, and their attachments of the post and present it to the LLM. The Discord to LLM connection code was based on llmcord [16] but was heavily adapted to our needs. The response from the LLM is added as a message to the post. All PETSc developers can see the message so that anyone can chime in. The message has three buttons: send, discard, and revise. If the response is fine, the developer clicks the send button to send the response to petsc-users with a signature of the name of the developer who clicked the button. [^2] The chatbot is still in private testing, and we haven’t sent any LLM-generated responses to the petsc-users mailing list. Also, the message in Discord will be time-stamped and tagged with the developer’s name. If the response is nonsense, developers can click the discard button to delete the message. Otherwise, if the response is partially correct and a developer clicks the revise button, the chatbot will ask the developer to guide a follow-up message. Developers can chat with the bot until a satisfying response is generated and ready to be sent. The Javascript for the Gmail account ignores emails from the chatbot and marks them as read, so that it doesn’t repost messages in the channel. In addition, users can do “direct messages” with the chatbot, keeping their conversation private; however, this may expose the user to unvetted hallucinations of the LLM.
@@
The above workflow leverages LLMs to help but not replace PETSc developers answering user questions from the mailing list. The negative example [10] alerts us not to let misleading or hallucinated answers damage PETSc’s reputation in the community.
V Preliminary Evaluation and Insights
V-A Evaluation strategy
@@
To evaluate the quality of answers, we leverage human expertise to classify the LLM responses using a blind-review workflow. Reviewers assigned a single score for each answer, based on correctness and completeness, without knowledge of the process used to generate it. The scoring rubric is presented in Table I. We compiled a benchmark dataset comprising 37 questions on the use of Krylov methods within PETSc, widely used by diverse scientific applications as a key part of composable algebraic solvers [17] and representative of a subtopic within the broader PETSc library.
@@
1: Incorrect or inaccurate statements (hallucinations) in the answer
@@
TABLE I: Rubric for LLM responses (higher is better).
V-B Comparison with LLM baseline

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: 54th International Conference on Parallel Processing Companion; September 08–11, 2025; San Diego, CA, USAbooktitle: 54th International Conference on Parallel Processing Companion (ICPP Companion ’25), September 08–11, 2025, San Diego, CA, USAdoi: 10.1145/3750720.3757281isbn: 979-8-4007-2109-0/25/09ccs: Mathematics of computing Solversccs: Information systems Language modelsccs: Software and its engineering Maintaining software
1. Introduction
High-performance scientific computing—which underpins discovery and innovation in all areas of science, engineering, technology, and society—relies on complex, evolving software libraries and tools that demand deep domain knowledge to develop, use, and maintain across continual advances in science drivers and computing architectures. Generative AI provides unprecedented opportunities to transform strategies for research and development (R&D) of scientific software, with the ultimate goal of accelerating scientific insight (Carter et al., 2023; Heroux et al., 2023; Papka et al., 2025). PETSc—the Portable, Extensible Toolkit for Scientific Computation—has served the scientific computing community for over three decades as a foundational software infrastructure for scalable numerical solvers (Balay et al., 2025; Mills et al., 2025; Adams et al., 2022). During this time, the project has cultivated a sophisticated and robust code base and a deep, distributed body of developer and user knowledge. The community knowledge base includes source code and documentation, email exchanges in public mailing lists, GitLab issues, git merge request reviews, Discord chats, manual pages and examples from the PETSc website, written materials such as papers and presentations, and importantly, PETSc developers and users’ understanding of PETSc. Among the written material, the structured and curated subset constitutes the official knowledge base: documentation, code, frequently asked questions (FAQs), and other resources that have undergone refinement and review, essentially everything that has passed a git merge request, while the remaining is the unofficial knowledge base. We call the subset of community knowledge that is available only in a developer’s brain (i.e., not written in a machine-accessible format) the wet knowledge base [^1] https://en.wikipedia.org/wiki/Wetware_(brain), and the opposite the dry knowledge base, see Fig. 1.
Figure 1. PETSc community knowledge base. AI assistants could help the flow of knowledge in all directions.
@@ 1. Introduction
These tools do not replace human contributors but act as scalable collaborators—suggesting code, summarizing discussions, generating documentation updates, and responding to user questions while adhering to PETSc’s collaborative development process. These same tools and techniques can and will be used to transfer material from the unofficial knowledge base to the official knowledge base, for example, dramatically improving manual pages with little developer effort. Developer-user (as well as developer-developer) communication is a common way that information is transferred from the wet knowledge base to machine-accessible format. Questions motivate developers to write down (often informally) knowledge and conceptualizations that they have never written down before. By having the LLM process these developer-user interactions, we can effectively capture wet knowledge as the developers perform their daily tasks without requiring them to try to core-dump their wet knowledge actively.
@@
Figure 2. Our goal is to create PETSc AI assistants capable of understanding and executing scientific computing tasks at the heart of numerical library research, providing the AI equivalent of human PETSc developers using LLM technology, while also continually advancing the PETSc knowledge base.
@@
Assistants embedded in tools such as Discord can answer user queries, recommend solutions, aid onboarding around PETSc’s broad functionality, and even analyze recurring feedback to guide improvements. For PETSc, these capabilities help address questions on the development of applications that build on PETSc scalable linear, nonlinear, and timestepping solvers as well as lower-level vector and matrix data structures for heterogeneous computing architectures (Zhang et al., 2022) and outer-loop numerical optimization and analytics.
@@
AI assistants can help in code review, suggest code improvements, aid in API design, offer performance tuning strategies, assist with refactoring complex code, and develop completely new functionality. Such assistants could help with continual advances needed in the PETSc code base to achieve efficient performance and address next-generation science challenges on ever-changing computer architectures.
@@
LLMs can help generate automated tests, detect bugs, and assist in debugging or improving continuous integration practices. Such capabilities would help open-source libraries like PETSc achieve comprehensive, multilevel testing to ensure a robust and reliable code base as part of the broader scientific software ecosystem.
@@
Beyond software development and maintenance, LLMs can support scientific discovery itself—by aiding in literature reviews (Antu et al., 2023), surfacing unexplored research directions (Schmidgall et al., 2025), and suggesting algorithms for novel problems.
@@
Through meeting summarization, task prioritization, and resource tracking, AI tools can reduce cognitive load and coordination overhead among contributors. Such capabilities help with communication and synchronization across the multi-institutional distributed PETSc developer team.
@@
2. PETSc User Support
User support, which we have consistently provided for over thirty years, is a fundamental responsibility of the PETSc team—as it is for any open-source scientific software project—to disseminate knowledge and realize the full impact of the software. Consequently, user support is our first focus when exploiting the PETSc knowledge base with AI assistants. Like many scientific computing libraries, PETSc has been funded primarily by the U.S. government. However, not only has this funding been limited, but it also has largely focused on supporting research in numerical computing, leaving few resources available for software maintenance and user support. Over the years, PETSc has developed a large and diverse user base, with tens of thousands of users across dozens of scientific domains. As a result, user support has become a significant undertaking. For instance, in May 2025 alone, there were over 300 messages across PETSc mailing lists, Discord channels, and GitLab issues. While PETSc developers have remained deeply committed to supporting users, the volume and complexity of user inquiries can often overwhelm a small team, particularly when resolving issues that involve extended, multiday, or multiweek discussions. Thus, to ensure that this process remains sustainable in the long term, we are particularly interested in leveraging recent advances in AI to assist with user support.
Although PETSc users can employ general-purpose AI tools such as ChatGPT, these tools can produce incorrect or hallucinated responses when answering technical questions about PETSc. The reason is primarily that mainstream LLMs do not have access to sufficient PETSc information. For example, in January 2025, we posed a PETSc-specific question to ChatGPT-4o and observed such limitations firsthand.
@@ 2. PETSc User Support
We scored this response as 0: it was a hallucination to a question about “KSPBurb” – a fictitious term that follows the naming convention of PETSc Krylov solvers (see Table 1 for the rubric used in the assessment). Although later versions of ChatGPT gave better results, this example shows the risk of relying solely on mainstream LLMs. The example also reminds us of the dangers of incorporating AI into user support. As noted in (Ycombinator, 2025), a prominent platform experienced backlash when its AI agents, masquerading as human support staff, fabricated a nonexistent login policy in response to user inquiries that led to widespread account cancellations. To safeguard the integrity and reputation of PETSc, we initially have adopted a cautious approach, and to date we have avoided full automation in user support interactions.
@@
To address the above challenges, we propose in Section 3 an augmented PETSc LLM workflow that can effectively process and deliver PETSc-centric information, and in Section 4 we present an integrated LLMs and PETSc Discord bots system that can reduce the risk of hallucinated responses reaching users.
3. The PETSc Augmented LLM Workflow
We have developed a preliminary PETSc LLM workflow, illustrated in Fig. 3, which can be used in multiple settings. For example, we could use the LLM to answer questions from the PETSc Discord server, from mailing lists, or from a chat box on the PETSc homepage. For developers, we could even provide command line tools and integrated development environment (IDE) extensions to facilitate various use cases. Our approach builds on standard LLM tools, such as LangChain (Langchain.com, 2025), while incorporating practical adaptations for PETSc. Specifically, we do the following:
@@ 3. The PETSc Augmented LLM Workflow
Figure 3. Our approach for creating PETSc AI assistants builds on standard LLM ecosystem tools. The boxes 1, 2, 3, and 4 in the top bar indicate the sequence of processing for PETSc-specific LLM interactions.
3.1. Generating the RAG databases
Retrieval-augmented generation algorithms ensure that users receive the most contextually appropriate and accurate information. RAG consists of two phases: the generation phase and the retrieval phase. Here, we discuss the generation phase. PETSc documentation is maintained in Markdown format and processed via Sphinx (The Sphinx Developers, 2024) into the website format. To construct a practical RAG database using this documentation, we use the following steps (see Fig. 3, beginning with “PETSc knowledge base”, following the blue arrows to “Embedding generator” and then the green arrow to “RAG databases”):
@@ 3.1. Generating the RAG databases
3.2. Utilizing the documentation RAG database
@@
3.3. Augmenting RAG searches with PETSc-specific keyword searches
@@ 3.4. Reranking enhanced RAG
Figure 4. Reranking-enhanced RAG workflow.
3.5. Postprocessing LLM output for users
@@
3.6. Storing and tracking interaction history
In addition to visually presenting the LLM output, we keep a detailed, manipulatable, searchable database of all interactions with all the LLMs by the PETSc developers (and eventually also users). In Fig. 3 this is depicted with the dotted orange arrow from box 4 to “Shared histories”; the blue dotted line to box 1 indicates that material from the shared history will also eventually be included in the RAG and reranking processing and passed to the LLM. Questions, responses, dates when done, and the continuation and embedding model used for each interaction will be included, as well as the generated prompts in the interaction. We provide an easy way for “scorers” to “blind-score” the responses to questions and indicate correct and incorrect portions of the responses. We can also score answers from PETSc developers stored in the same database, establishing high quality and trust. Currently, we use a bespoke Python dictionary to manage the database, but we are pursuing using emerging agentic memory systems (Hatalis et al., 2023).
4. PETSc AI Bots in Discord
Discord (Discord, 2025) is an instant messaging platform where communication happens in virtual communities called “servers.” We have set up a PETSc Discord server, and hundreds of users have joined since its debut. We integrated augmented PETSc LLMs with the Discord server and set up a chatbot that has two usage modes: (1) one can directly chat with it, or (2) the chatbot can answer questions from emails, while PETSc developers can monitor and vet answers on back channels. In this section we first briefly introduce Discord’s integration mechanisms, then describe the workflow that connects a PETSc public mailing list and LLMs; see Fig. 5.
Figure 5. Bots in Discord that connect a PETSc mailing list and an LLM. Numbered arcs show a typical sequence of events in the workflow.
@@ 4. PETSc AI Bots in Discord
PETSc has three public mailing lists: {petsc-users, petsc-maint, petsc-dev}. petsc-dev is mainly for PETSc developers and others interested in PETSc development. The other two are for PETSc users with one distinction: petsc-maint is a private maintenance email with controlled recipients and no public archives, while petsc-users is fully public and has archives from the past twenty years. In this study we targeted petsc-users but didn’t touch its archives for RAG. We created a Google Gmail account petscbot@gmail.com and subscribed it to the petsc-users mailing list. Then, with Google Apps Script services (Google, 2025), we use JavaScript to periodically check whether there are new (unread) emails from petsc-users in the Gmail account. If there are, the script sends a message to a webhook associated with a private channel named petsc-users-notification in PETSc Discord. An email bot (application) was added to the PETSc Discord server, which watches for new messages in petsc-users-notification. Whenever there is a new message, using Gmail’s Python API, the email bot fetches unread emails from petsc-users in the Gmail account and marks them read. The fetched emails are posted to a private channel, visible only to PETSc developers, named petsc-users-emails. This channel is a Discord Forum channel made of posts. We can make each email thread a post, with its title being the email thread’s subject and emails within the thread follow up messages in the post. We also fetch and attach email attachments to their corresponding messages in Discord. We lightly parse email bodies to remove quotes commonly seen in email replies and revert the url-defense protected URLs so that messages are presented concisely.
@@
From a post in petsc-users-emails, any PETSc developer can invoke a command /reply on the chatbot, which will then build a conversation context with the title, messages, and their attachments of the post and present it to the LLM. The Discord to LLM connection code was based on llmcord (Jakobdylanc, 2025) but was heavily adapted to our needs. The response from the LLM is added as a message to the post. All PETSc developers can see the message so that anyone can chime in. The message has three buttons: send, discard, and revise. If the response is fine, the developer clicks the send button to send the response to petsc-users with a signature of the name of the developer who clicked the button. [^2] The chatbot is still in private testing, and we haven’t sent any LLM-generated responses to the petsc-users mailing list. Also, the message in Discord will be time-stamped and tagged with the developer’s name. If the response is nonsense, developers can click the discard button to delete the message. Otherwise, if the response is partially correct and a developer clicks the revise button, the chatbot will ask the developer to guide a follow-up message. Developers can chat with the bot until a satisfying response is generated and ready to be sent. The Javascript for the Gmail account ignores emails from the chatbot and marks them as read, so that it doesn’t repost messages in the channel. In addition, users can do “direct messages” with the chatbot, keeping their conversation private; however, this may expose the user to unvetted hallucinations of the LLM.
@@
The above workflow leverages LLMs to help but not replace PETSc developers answering user questions from the mailing list. The negative example (Ycombinator, 2025) alerts us not to let misleading or hallucinated answers damage PETSc’s reputation in the community.
5. Preliminary Evaluation and Insights
5.1. Evaluation strategy
@@
To evaluate the quality of answers at the highest standard, we leverage human expertise to classify the LLM responses using a blind-review workflow. Two experienced PETSc developers (with 10+ years of experience) served as the reviewers; they assigned a single score for each answer, based on correctness and completeness, without knowledge of the process used to generate it. The scoring rubric is presented in Table 1. We compiled a benchmark dataset comprising 37 questions on the use of Krylov methods within PETSc, widely used by diverse scientific applications as a key part of composable algebraic solvers (Brown et al., 2012) and representative of a subtopic within the broader PETSc library. The questions were derived from real user inquiries and span a range of difficulty levels, including both easy and hard questions. The hard questions are often open-ended and cannot be answered directly using the PETSc documentation.
@@
1: Incorrect or inaccurate statements (hallucinations)
@@
Table 1. Rubric for LLM responses (higher is better).
5.2. Comparison with LLM baseline